# Prediction-Powered Adaptive Shrinkage Estimation

### v2

@@ 4.1 The Within Problem Power-Tuning Stage
The formulation of the above PT estimators is well understood in the single problem setting (Angelopoulos et al., 2024; Miao et al., 2024a). In PAS, we execute this stage separately for each problem, as the optimal power-tuning parameter is problem-dependent and varies case by case.
@@ 4.2 The Across Problem Adaptive Shrinkage Stage
$\hat{\theta}_{j,\omega}^{\mathrm{PAS}}$ $:=\omega_{j}\hat{\theta}^{\mathrm{PT}}_{j}+(1-\omega_{j})\tilde{Z}^{f}_{j},$ (15)
$\text{with}\quad\omega_{j}$ $\equiv\omega_{j}(\omega):=\frac{\omega}{\omega+\tilde{\sigma}^{2}_{j}}.$
@@
then the resulting estimator $\hat{\boldsymbol{\theta}}_{\omega^{*}}^{\mathrm{PAS}}$ would satisfy all our desiderata. While this construction is not feasible since the compound risk function in (9) depends on the unknown $\boldsymbol{\eta},\boldsymbol{\theta}$, we can make progress by pursuing a classical statistical idea: we can develop an unbiased estimate of the compound risk (Mallows, 1973; Stein, 1981; Efron, 2004) and then use it as a surrogate for tuning $\omega$.
@@
See Appendices B and F.1 for the proof and motivation. With Theorem 4.1 in hand, we now have a systematic strategy of picking $\omega$ by minimizing CURE, following the paradigm of tuning parameter selection via minimization of an unbiased risk estimate (as advocated by, e.g. Li (1985); Donoho & Johnstone (1995); Xie et al. (2012); Candès et al. (2013); Ignatiadis & Wager (2019); Ghosh et al. (2025)): [^11] The connection to EB is the following. Xie et al. (2012) and Tibshirani & Rosset (2019) explain that James-Stein-type estimators may be derived by tuning $\sigma_{\theta}^{2}$ (in Section 3) via minimization of Stein’s (1981) unbiased risk estimate (SURE).
@@
7:    $\tilde{\sigma}_{j}^{2}=\texttt{{get\_pt\_var}}(\hat{\theta}_{j}^{\mathrm{PT}})$triangleright (Eq. (14))
@@ 5 Asymptotic Optimality
Suppose the datasets are generated according to Assumptions 2.1 and 2.2 and further assume that $\>\mathbb{E}_{\mathbb{P}_{\eta}}[Y_{ij}^{4}]<\infty,\>\mathbb{E}_{\mathbb{P}_{\eta}}[f(X_{ij})^{4}]<\infty$. Then,
@@
where $o(1)$ denotes a term that converges to 00 as $m\to\infty$.
@@
$\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\Big{(}\boldsymbol{\hat{\theta}}^{\mathrm{PAS}}_{\hat{\omega}}\Big{)}\leq\inf_{\omega\geq 0}\left\{\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\Big{(}\boldsymbol{\hat{\theta}}^{\mathrm{PAS}}_{\omega}\Big{)}\right\}+o(1)\,\text{ as }\,m\to\infty,$
@@ 7 Experiments
    the vanilla PPI estimator (Angelopoulos et al., 2023);
@@
We include the estimators in (vii.) only for the real-world experiments, as they are specifically designed for settings where the second moments are unknown. See Appendix D for detailed formulations and implementations for (v.) & (vi.) (the latter of which is inspired by the SURE-grand mean estimator of Xie et al. (2012)). The comparisons with (iv.) and (v.) also directly serve as ablation studies of the two stages in constructing the PAS estimator.
@@
For real-world datasets, we introduce a second metric to assess whether improvements in MSE are driven by a few difficult problems rather than consistent performance gains: the percentage of problems improved relative to the classical estimator (abbreviated as “$\mathrm{Improved}\,\%\,\uparrow$”). This metric is defined as $\frac{1}{m}\sum_{j=1}^{m}\mathbf{1}[(\hat{\theta}_{j}-\theta_{j})^{2}<(\hat{\theta}_{j}^{\mathrm{Classical}}-\theta_{j})^{2}]\times(100\%).$ Larger values of this metric are preferable.
@@ 7.1 Synthetic Model
Estimator MSE $f_{1}$ (×$\times 10^{-3}$) MSE $f_{2}$ (×$\times 10^{-3}$)
@@
Shrink Classical 0.273 $\pm$ 0.003 3.817 $\pm$ 0.042
Shrink Avg 2.486 $\pm$ 0.026 2.575 $\pm$ 0.026
PAS (ours) 0.272 $\pm$ 0.003 2.496 $\pm$ 0.025
@@
For both predictors, we see that PAS outperforms all the baselines. With a good predictor $f_{1}$, both the prediction mean and the shrinkage estimator closely track PAS; in contrast, the PPI and PT estimators fail to fully leverage the accurate predictions, as their design enforces unbiasedness. The situation reverses for the less reliable predictor $f_{2}$: the prediction mean and the shrinkage estimator have high MSE, while estimators with built-in debiasing mechanisms demonstrate greater resilience. PAS adapts effectively across these extremes, making it a handy choice for a wide range of problems and predictors.
@@
Amazon (base $f$) Amazon (tuned $f$) Galaxy Estimator MSE (×$(\times 10^{-3})$ Improved $\uparrow$ MSE (×$(\times 10^{-3})$ Improved $\uparrow$ MSE (×$(\times 10^{-3})$ Improved $\uparrow$ Classical 24.305 $\pm$ 0.189 baseline 24.305 $\pm$ 0.189 baseline 2.073 $\pm$ 0.028 baseline Prediction Avg 41.332 $\pm$ 0.050 30.7 $\pm$ 0.2 3.945 $\pm$ 0.011 75.4 $\pm$ 0.2 7.195 $\pm$ 0.008 17.0 $\pm$ 0.2 PPI 11.063 $\pm$ 0.085 62.4 $\pm$ 0.2 7.565 $\pm$ 0.066 70.4 $\pm$ 0.2 1.149 $\pm$ 0.017 59.4 $\pm$ 0.3 PT 10.633 $\pm$ 0.089 70.3 $\pm$ 0.2 6.289 $\pm$ 0.050 76.0 $\pm$ 0.2 1.026 $\pm$ 0.015 67.7 $\pm$ 0.3 Shrink Classical 15.995 $\pm$ 0.121 56.4 $\pm$ 0.3 3.828 $\pm$ 0.039 78.9 $\pm$ 0.2 1.522 $\pm$ 0.016 48.8 $\pm$ 0.4 Shrink Avg 9.276 $\pm$ 0.078 70.4 $\pm$ 0.2 6.280 $\pm$ 0.058 77.1 $\pm$ 0.2 0.976 $\pm$ 0.014 68.9 $\pm$ 0.3 PAS (ours) 8.517 $\pm$ 0.071 71.4 $\pm$ 0.2 3.287 $\pm$ 0.024 80.8 $\pm$ 0.2 0.893 $\pm$ 0.011 67.3 $\pm$ 0.4 UniPT (ours) 10.272 $\pm$ 0.084 70.0 $\pm$ 0.2 6.489 $\pm$ 0.053 76.2 $\pm$ 0.2 1.017 $\pm$ 0.015 67.7 $\pm$ 0.3 UniPAS (ours) 8.879 $\pm$ 0.073 69.5 $\pm$ 0.2 3.356 $\pm$ 0.031 77.6 $\pm$ 0.2 0.909 $\pm$ 0.011 66.7 $\pm$ 0.3
@@ 7.2 Real-World Datasets
Amazon Review Ratings (SNAP, 2014).   Many commercial and scientific studies involve collecting a large corpus of text and estimating an average score (rating, polarity, etc.) from it. A practitioner would often combine limited expert annotations with massive automatic evaluations from ML models (Baly et al., 2020; Egami et al., 2023; Fan et al., 2024; Mozer & Miratrix, 2025). To emulate this setup, we consider mean rating estimation problems using the Amazon Fine Food Review dataset from Kaggle, where we artificially hide the labels in a random subset of the full data to serve as the unlabeled partition. Concretely, we estimate the average rating for the top $m=200$ products with the most reviews (from $\sim$200 to $\sim$900). For the $i$-th review of the $j$-th product, the covariate $X_{ij}$ consists of the review’s title and text concatenated, while the outcome $Y_{ij}$ is the star rating in $\{1,\dots,5\}$. We employ two black-box predictors: (1) BERT-base, a language model without fine-tuning (Devlin et al., 2019) and (2) BERT-tuned which is the same model but fine-tuned on a held-out set of reviews from other products. Neither of them are trained on the reviews of the 200 products we evaluate.
Spiral Galaxy Fractions (Willett et al., 2013).   The Galaxy Zoo 2 project contains the classification results of galaxy images from the the Sloan Digital Sky Survey (York et al., 2000, SDSS). We are interested in estimating the fraction of galaxies that are classified as “spiral,” i.e., have at least one spiral arm. The covariates $X_{ij}$ in this applications are images (we provide some examples in Figure 5 of the appendix). Existing PPI papers have focused on estimating the overall fraction; we demonstrate how this dataset’s metadata structure enables compound estimation of spiral fractions across distinct galaxy subgroups. We first use a pre-defined partition of the galaxies into 100 subgroups that is based on the metadata attribute WVT_BIN. Second, we estimate the fraction of spiral galaxies in all of the galaxy subgroups simultaneously. The predictor is a ResNet50 network trained on a held-out set with around 50k images.
@@ 8 Conclusion
This paper introduces PAS, a novel method for compound mean estimation that effectively combines PPI and EB principles. We motivate the problem through the lens of variance reduction and contextual prior information—then demonstrate how PAS achieves both goals, in theory and in practice. Our paper differs from many other PPI-related works in its focus on estimation, so a natural next step is to develop average coverage controlling intervals for the means centered around PAS. To this end, it may be fruitful to build on the robust empirical Bayes confidence intervals of Armstrong et al. (2022). Modern scientific inquiries increasingly demand the simultaneous analysis of multiple related problems. The framework developed in this paper represents a promising direction for such settings.
@@ A.1 Further Connections to Related Works
Stratified PPI (StratPPI; Fisch et al. (2024)).
Fisch et al. (2024) also consider a setting with stratification of observations into subproblems. However, their overall aim is still to estimate a single parameter, rather than multiple parameters, as we do. Thus, in solving individual problems, for instance, they still adhere to the requirement of unbiasedness. As an example, let us revisit the Galaxy Zoo 2 application. Therein, we mentioned the following distinction between our work and previous methods in the PPI literature:
  • Estimate the fraction $\theta$ of spiral galaxies across the whole universe (Angelopoulos et al., 2023).
@@
$(X,Y)\,{\buildrel\text{iid}\over{\sim}\,}\mathbb{P}.$
@@
PPI++ (Angelopoulos et al., 2024) with multivariate mean estimand.
Angelopoulos et al. (2024) (and other papers in the PPI literature) consider statistical problems going beyond the estimation of a univariate mean. For instance, the results of Angelopoulos et al. (2024) also accommodate multivariate mean estimands. Here we explain how one can frame our compound mean estimation setting into the estimation of a single but multivariate mean estimand, which then fits in the more generalized formulation for power tuning considered in PPI++.
@@
By independence across problems, we have the covariance structures $\mathrm{Var}[f(\tilde{X}_{i.})]=\text{diag}((\tau_{j}^{2})_{j=1}^{m})$ and $\mathrm{Cov}[f(X_{i.}),Y_{i.}]=\text{diag}((\gamma_{j})_{j=1}^{m})$ where $\text{diag}(\cdot)$ constructs a diagonal matrix with its arguments on the diagonal. Angelopoulos et al. (2024) then consider a class of estimators indexed by a single weighting parameter $\lambda$ that is applied to the entire vector:
@@ A.2.1 Why Does Sharing Information Across Problems Help?
Here we provide an explanation of the fundamental statistical difference between estimating a single mean versus estimating a lot of means ($m\to\infty$) with an eye toward providing intuition of how we can share information across problems. We emphasize that this is only meant for intuition; the empirical Bayes literature, starting with James & Stein (1961), provides a more nuanced picture.
@@ A.2.2 Analogy between $Z_{j}^{f}$ and $\mathbb{E}[\theta\mid\psi]$
[…]
@@
    The above expression is interesting as it forces us to inflate “$\omega$”, i.e., to shrink less toward $h(\phi)$ in a way that depends on how close $h(\phi)$ is to $\mathbb{E}[\theta\mid\phi]$. See Ignatiadis & Wager (2019, last paragraph of Section 3) and Section A.2.3 below for further discussion of this point.

### v3

@@ 4.1 The Within Problem Power-Tuning Stage
The formulation of the above PT estimators is well understood in the single problem setting (angelopoulos_ppi_2024; miao_assumption-lean_2024). In PAS, we execute this stage separately for each problem, as the optimal power-tuning parameter is problem-dependent and varies case by case.
@@ 4.2 The Across Problem Adaptive Shrinkage Stage
$\hat{\theta}_{j,\omega}^{\mathrm{PAS}}$ $=\omega_{j}\hat{\theta}^{\mathrm{PT}}_{j}+(1-\omega_{j})\tilde{Z}^{f}_{j},$ (15)
$\text{with}\quad\omega_{j}$ $\equiv\omega_{j}(\omega)=\frac{\omega}{\omega+\tilde{\sigma}^{2}_{j}}.$
@@
then the resulting estimator $\hat{\boldsymbol{\theta}}_{\omega^{*}}^{\mathrm{PAS}}$ would satisfy all our desiderata. While this construction is not feasible since the compound risk function in (9) depends on the unknown $\boldsymbol{\eta},\boldsymbol{\theta}$, we can make progress by pursuing a classical statistical idea: we can develop an unbiased estimate of the compound risk (mallows1973comments; stein1981estimation; efron2004estimation) and then use it as a surrogate for tuning $\omega$.
@@
See Appendices B and F.1 for the proof and motivation. With Theorem 4.1 in hand, we now have a systematic strategy of picking $\omega$ by minimizing CURE, following the paradigm of tuning parameter selection via minimization of an unbiased risk estimate (as advocated by, e.g. li1985stein; donoho1995adapting; xie2012sure; candes2013unbiased; ignatiadis2019covariatepowered; ghosh2025steins): [^11] The connection to EB is the following. xie2012sure and tibshirani2019excess explain that James-Stein-type estimators may be derived by tuning $\sigma_{\theta}^{2}$ (in Section 3) via minimization of Stein’s (stein1981estimation) unbiased risk estimate (SURE).
@@
7:  $\tilde{\sigma}_{j}^{2}=\texttt{{get\_pt\_var}}(\hat{\theta}_{j}^{\mathrm{PT}})$ $\triangleright$ (Eq. (14))
@@ 5 Asymptotic Optimality
Suppose the datasets are generated according to Assumptions 2.1 and 2.2 and further assume that $\mathbb{E}_{\mathbb{P}_{\eta}}[Y_{ij}^{4}]<\infty,\>\mathbb{E}_{\mathbb{P}_{\eta}}[f(X_{ij})^{4}]<\infty$. Then,
@@
where $o(1)$ denotes a term that converges to 0 as $m\to\infty$.
@@
[…]
@@ 7 Experiments
    the vanilla PPI estimator (angelopoulos2023prediction);
@@
We include the estimators in (vii.) only for the real-world experiments, as they are specifically designed for settings where the second moments are unknown. See Appendix D for detailed formulations and implementations for (v.) & (vi.) (the latter of which is inspired by the SURE-grand mean estimator of xie2012sure). The comparisons with (iv.) and (v.) also directly serve as ablation studies of the two stages in constructing the PAS estimator.
@@
For real-world datasets, we introduce a second metric to assess whether improvements in MSE are driven by a few difficult problems rather than consistent performance gains: the percentage of problems improved relative to the classical estimator (abbreviated as $\mathrm{Improved}\,\%\,\uparrow$”). This metric is defined as $\frac{1}{m}\sum_{j=1}^{m}\mathbf{1}[(\hat{\theta}_{j}-\theta_{j})^{2}<(\hat{\theta}_{j}^{\mathrm{Classical}}-\theta_{j})^{2}]\times(100\%).$ Larger values of this metric are preferable.
@@ 7.1 Synthetic Model
Estimator MSE $f_{1}$ ($\times 10^{-3}$) MSE $f_{2}$ ($\times 10^{-3}$)
@@
Shrink Classical 0.272 $\pm$ 0.003 2.863 $\pm$ 0.030
Shrink Avg 2.486 $\pm$ 0.026 2.537 $\pm$ 0.026
PAS (ours) 0.272 $\pm$ 0.003 2.466 $\pm$ 0.026
@@
For both predictors, we see that PAS is (one of) the best among all estimators. With a good predictor $f_{1}$, both the prediction mean and the shrinkage estimator closely track PAS; in contrast, the PPI and PT estimators fail to fully leverage the accurate predictions, as their design enforces unbiasedness. The situation reverses for the less reliable predictor $f_{2}$: the prediction mean and the shrinkage estimator have high MSE, while estimators with built-in debiasing mechanisms demonstrate greater resilience. PAS adapts effectively across these extremes, making it a handy choice for a wide range of problems and predictors.
@@
Amazon (base $f$) Amazon (tuned $f$) Galaxy Estimator MSE $(\times 10^{-3})$ % Improved $\uparrow$ MSE $(\times 10^{-3})$ % Improved $\uparrow$ MSE $(\times 10^{-3})$ % Improved $\uparrow$ Classical 24.305 $\pm$ 0.189 baseline 24.305 $\pm$ 0.189 baseline 2.073 $\pm$ 0.028 baseline Prediction Avg 41.332 $\pm$ 0.050 30.7 $\pm$ 0.2 3.945 $\pm$ 0.011 75.4 $\pm$ 0.2 7.195 $\pm$ 0.008 17.0 $\pm$ 0.2 PPI 11.063 $\pm$ 0.085 62.4 $\pm$ 0.2 7.565 $\pm$ 0.066 70.4 $\pm$ 0.2 1.149 $\pm$ 0.017 59.4 $\pm$ 0.3 PT 10.633 $\pm$ 0.089 70.3 $\pm$ 0.2 6.289 $\pm$ 0.050 76.0 $\pm$ 0.2 1.026 $\pm$ 0.015 67.7 $\pm$ 0.3 Shrink Classical 15.995 $\pm$ 0.121 56.4 $\pm$ 0.3 3.828 $\pm$ 0.039 78.9 $\pm$ 0.2 1.522 $\pm$ 0.016 48.8 $\pm$ 0.4 Shrink Avg 9.276 $\pm$ 0.078 70.4 $\pm$ 0.2 6.280 $\pm$ 0.058 77.1 $\pm$ 0.2 0.976 $\pm$ 0.014 68.9 $\pm$ 0.3 PAS (ours) 8.517 $\pm$ 0.071 71.4 $\pm$ 0.2 3.287 $\pm$ 0.024 80.8 $\pm$ 0.2 0.893 $\pm$ 0.011 67.3 $\pm$ 0.4 UniPT (ours) 10.272 $\pm$ 0.084 70.0 $\pm$ 0.2 6.489 $\pm$ 0.053 76.2 $\pm$ 0.2 1.017 $\pm$ 0.015 67.7 $\pm$ 0.3 UniPAS (ours) 8.879 $\pm$ 0.073 69.5 $\pm$ 0.2 3.356 $\pm$ 0.031 77.6 $\pm$ 0.2 0.909 $\pm$ 0.011 66.7 $\pm$ 0.3
@@ 7.2 Real-World Datasets
Amazon Review Ratings (amazon_fine_food_reviews).   Many commercial and scientific studies involve collecting a large corpus of text and estimating an average score (rating, polarity, etc.) from it. A practitioner would often combine limited expert annotations with massive automatic evaluations from ML models (baly2020we; egami2023using; fan2024narratives; mozer2025more). To emulate this setup, we consider mean rating estimation problems using the Amazon Fine Food Review dataset from Kaggle, where we artificially hide the labels in a random subset of the full data to serve as the unlabeled partition. Concretely, we estimate the average rating for the top $m=200$ products with the most reviews (from $\sim$200 to $\sim$900). For the $i$-th review of the $j$-th product, the covariate $X_{ij}$ consists of the review’s title and text concatenated, while the outcome $Y_{ij}$ is the star rating in $\{1,\dots,5\}$. We employ two black-box predictors: (1) BERT-base, a language model without fine-tuning (devlin2019bert) and (2) BERT-tuned which is the same model but fine-tuned on a held-out set of reviews from other products. Neither of them are trained on the reviews of the 200 products we evaluate.
Spiral Galaxy Fractions (willett2013galaxy).   The Galaxy Zoo 2 project contains the classification results of galaxy images from the the Sloan Digital Sky Survey (york2000sloan, SDSS). We are interested in estimating the fraction of galaxies that are classified as “spiral,” i.e., have at least one spiral arm. The covariates $X_{ij}$ in this applications are images (we provide some examples in Figure 5 of the appendix). Existing PPI papers have focused on estimating the overall fraction; we demonstrate how this dataset’s metadata structure enables compound estimation of spiral fractions across distinct galaxy subgroups. We first use a pre-defined partition of the galaxies into 100 subgroups that is based on the metadata attribute WVT_BIN. Second, we estimate the fraction of spiral galaxies in all of the galaxy subgroups simultaneously. The predictor is a ResNet50 network trained on a held-out set with around 50k images.
@@ 8 Conclusion
This paper introduces PAS, a novel method for compound mean estimation that effectively combines PPI and EB principles. We motivate the problem through the lens of variance reduction and contextual prior information—then demonstrate how PAS achieves both goals, in theory and in practice. Our paper differs from many other PPI-related works in its focus on estimation, so a natural next step is to develop average coverage controlling intervals for the means centered around PAS. To this end, it may be fruitful to build on the robust empirical Bayes confidence intervals of armstrong2022robust. Modern scientific inquiries increasingly demand the simultaneous analysis of multiple related problems. The framework developed in this paper represents a promising direction for such settings.
@@ Acknowledgments
$\clubsuit$ Appendix: Table of Contents
  1. A.
    Further Connections and Intuitions Behind PAS ........................................................................................................................................................................Page A
  2. B.
    The Correlation-Aware Unbiased Risk Estimate ........................................................................................................................................................................Page B
  3. C.
    Details on PAS with Unknown Second Moments ........................................................................................................................................................................Page C
  4. D.
    Baseline Shrinkage Estimators ........................................................................................................................................................................Page D
  5. E.
    Experiment Details ........................................................................................................................................................................Page E
  6. F.
    Proofs of Theoretical Results ........................................................................................................................................................................Page F
@@ A.1 Further Connections to Related Works
Stratified PPI (StratPPI; fisch2024stratified).
fisch2024stratified also consider a setting with stratification of observations into subproblems. However, their overall aim is still to estimate a single parameter, rather than multiple parameters, as we do. Thus, in solving individual problems, for instance, they still adhere to the requirement of unbiasedness. As an example, let us revisit the Galaxy Zoo 2 application. Therein, we mentioned the following distinction between our work and previous methods in the PPI literature:
  • Estimate the fraction $\theta$ of spiral galaxies across the whole universe (angelopoulos2023prediction).
@@
$(X,Y)\,{\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\text{iid}}}\,}\mathbb{P}.$
@@
PPI++ (angelopoulos_ppi_2024) with multivariate mean estimand.
angelopoulos_ppi_2024 (and other papers in the PPI literature) consider statistical problems going beyond the estimation of a univariate mean. For instance, the results of angelopoulos_ppi_2024 also accommodate multivariate mean estimands. Here we explain how one can frame our compound mean estimation setting into the estimation of a single but multivariate mean estimand, which then fits in the more generalized formulation for power tuning considered in PPI++.
@@
By independence across problems, we have the covariance structures $\mathrm{Var}[f(\tilde{X}_{i.})]=\text{diag}((\tau_{j}^{2})_{j=1}^{m})$ and $\mathrm{Cov}[f(X_{i.}),Y_{i.}]=\text{diag}((\gamma_{j})_{j=1}^{m})$ where $\text{diag}(\cdot)$ constructs a diagonal matrix with its arguments on the diagonal. angelopoulos_ppi_2024 then consider a class of estimators indexed by a single weighting parameter $\lambda$ that is applied to the entire vector:
@@ A.2.1 Why Does Sharing Information Across Problems Help?
Here we provide an explanation of the fundamental statistical difference between estimating a single mean versus estimating a lot of means ($m\to\infty$) with an eye toward providing intuition of how we can share information across problems. We emphasize that this is only meant for intuition; the empirical Bayes literature, starting with james1961estimation, provides a more nuanced picture.
@@ A.2.2 Analogy between $Z_{j}^{f}$ and $\mathbb{E}[\theta\mid\psi]$
for some $\omega\geq 0$. [^16] At this stage, there is 1-1 mapping between $\omega\geq 0$ and $w\in[0,1]$. However, for PAS, $\sigma_{\varepsilon}^{2}$ corresponds to the variance of the $j$-th power tuning estimator and will be different from problem to problem. Then it will be convenient to parameterize all the problems by the same parameter $\omega\geq 0$.
@@
    The above expression is interesting as it forces us to inflate $\omega$”, i.e., to shrink less toward $h(\phi)$ in a way that depends on how close $h(\phi)$ is to $\mathbb{E}[\theta\mid\phi]$. See ignatiadis2019covariatepowered and Section A.2.3 below for further discussion of this point.